OpenStudy (anonymous):

Solve: –4(2x – 1) = –6(x + 2) – 2

OpenStudy (anonymous):

-4(2x-1) = -6(x+2) -2 -8x+4 = -6x - 12 - 2 -8x+4 = -6x - 14 4 = 2x -14 2x = 18 x = 9

OpenStudy (anonymous):

(-2)-1..?

OpenStudy (anonymous):

Solving it so.. Shouldn't you be solving for x, which would be x = 9 when you seperate it from the other numbers?

OpenStudy (anonymous):

i got 9 too i was just checking if i did it right... thanks and your solving for x.